OpenStudy (iwanttogotostanford):

please help:

OpenStudy (iwanttogotostanford):

@phi @TheSmartOne @sweetburger @pooja195

OpenStudy (phi):

I would change tangent into sin/cos and do the same for cotangent. Do you know how ?

OpenStudy (iwanttogotostanford):

i got tan^2 theta

OpenStudy (iwanttogotostanford):

@phi

OpenStudy (phi):

yes, that looks ok \[ \frac{ \frac{\sin x }{\cos x}}{ \frac{\cos x}{\sin x }} \\ \] multiply top and bottom by sinx / cosx you get \[ \frac{ \frac{\sin^2 x }{\cos^2 x}}{1}\] which is the same as \[ \frac{\sin^2 x }{\cos^2 x} = \left(\frac{\sin x }{\cos x} \right)^2 = \tan^2 x \]

OpenStudy (iwanttogotostanford):

thanks! @phi can I have help with one more please?

OpenStudy (phi):

did you make a new post ?

TheSmartOne (thesmartone):

It would have been easier if you used the fact that cot = 1/tan \(\Large \frac{tan\theta}{cot\theta} = tan\theta \div cot \theta = \) \(\Large tan\theta\div \frac{1}{tan \theta} = tan \theta \times tan \theta = \boxed{\bf{tan^2\theta}}\)